# arXiv:2603.01316v1[eess.AS] 01 Mar 2026

### v1

@@ IV-D Relative Cues versus Independent Cues
Experimental Setup: We evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework. For discrete-valued attributes, relative and independent cues provide the same level of discriminability; therefore, we restrict our evaluation to originally continuous-valued attributes. In total, eight cues are considered, of which two—age category and temporal order—are excluded. The age category cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable. Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (youth: 10,310; young: 7,063; middle: 2,627; senior: 0). The temporal order cue, defined as the difference in onset time between the target and interfering speech, is inherently relational and therefore does not admit a meaningful absolute categorization. Consequently, the evaluation focuses on the remaining six cues: pitch level, pitch range, loudness, distance, speaking rate, and speaking duration.
Separated speech samples are selected from the TF-Locoformer separation model corresponding to these cues. Two binary classification models are trained: one guided by relative-cue text prompts and the other by independent-cue text prompts.
TABLE VII: Metrics for Independent and Relative Cues
@@
Pitch Level non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
Pitch Level all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
@@
Independent Cues
  • Pitch level: low / normal / high
  • Pitch range: narrow / normal / wide
  • Loudness: quiet / normal / loud
  • Distance: near / far
  • Speaking rate: slow / normal / fast
  • Speaking duration: short / long
Relative Cues
Relative cue categories are defined by comparing the target speech with the interference speech:
  • Pitch level: lower / similar / higher
  • Pitch range: narrower / similar / wider
  • Loudness: quieter / similar / louder
  • Distance: nearer / similar / farther
  • Speaking rate: slower / similar / faster
  • Speaking duration: shorter / similar / longer
We use the audio data corresponding to six relative cues—pitch level, pitch range, loudness, distance, speaking rate, and speaking duration—from the dataset described in Section IV. To isolate the effect of cue category differences, we generate prompts for independent cues and regenerate prompts for relative cues using predefined templates. Each prompt describes a single cue, and all prompts share the same sentence structure, differing only in whether the cue is expressed in relative or independent terms. For example, a relative-cue prompt is “Please extract the speaker with a higher pitch level in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch level in the audio.”
During training, samples in which both speech signals belong to the same absolute category (or the Similar relative category) are excluded. During evaluation, all samples are retained, and target labels are derived from the PIT results of the separation model.
Classification accuracy is reported under different conditions. For independent cues, three conditions are considered: (i) same category, (ii) adjacent categories (e.g., slow–normal for speaking rate), and (iii) distinctly different categories (e.g., slow–fast for speaking rate). For relative cues, two conditions are considered: (i) same category (“Similar”) and (ii) different categories. We further analyze cases in which target–interference pairs fall into different relative-cue categories but are quantized into the same, adjacent, or distinct absolute categories, highlighting scenarios where independent cues provide limited discriminability while relative cues remain informative.
@@
Results and Analysis: Overall classification accuracy is compared across cue types to evaluate the effectiveness of relative versus independent cues for TSE. To facilitate analysis, pairs are grouped as follows: for independent cues, pairs in the same category are labeled the same group, and pairs in different categories are labeled the non-same group; for relative cues, pairs in the same category are labeled the similar group, and pairs in different categories are labeled the non-similar group. Results are presented across six continuous-valued attributes.
For independent cues, we report the classification accuracy, as well as the resulting SI-SDRi and PESQ performance for the non-same and same groups. For relative cues, we report the classification accuracy and the corresponding SI-SDRi and PESQ performance for the non-similar and similar groups. We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues. Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples. The results are summarized in Table VII.
The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues. Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information. In particular, for pitch level, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
For the distance and loudness cues, using relative categories results in more samples being assigned to the similar group than when using absolute categories. This is due to the limited ranges of the simulated speaker-to-microphone distances and SIR values. In our dataset, the speaker-to-microphone distances in the simulated mixtures range from 0.3 to 1.5 m, and an auditory distance threshold of 0.5 m is used to distinguish between the two speakers. Consequently, many target-interference speech pairs are classified as having similar distances. A similar effect is observed for the loudness cue: loudness is approximated using SIR, which is constrained to the range of -6 to 6 dB in our simulation. With a loudness cue threshold of 3 dB, nearly half of the samples are categorized as having similar loudness.
@@
It is informative to examine the specific categories of target–interference samples. Fig. 5 presents the classification accuracy and sample distribution for both relative and independent cues, using speaking rate as an example. When the target and interference speech fall into clearly distinct categories (e.g., slow–fast or fast–slow), the model achieves the highest accuracy. In contrast, accuracy decreases when they belong to adjacent categories (e.g., slow–normal or normal–fast), supporting the assumption in Section II.
To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as relative-different and examine whether they fall into the same, adjacent, or distinct absolute categories. Using speaking rate as an example (Fig. 6), we report classification accuracy under these three absolute conditions for both the full set and the relative-different subset, along with their sample distribution. Accuracy increases monotonically with the magnitude of speaking rate difference between two speakers. In the same condition, performance is relatively low (48.9%), but restricting to relative-different pairs improves accuracy to 79.4%, indicating that relative comparison provides additional discriminative information even when absolute categories are identical. In the adjacent condition, accuracy is high (87.3%) and slightly improves to 89.5% under the relative-different subset. In the distinct condition, performance reaches 98.3% in both cases. These results show that classification performance strongly correlates with speaking rate difference between two speakers, and that relative cues particularly benefit challenging cases where absolute quantization obscures pairwise distinctions.
@@ V Conclusion
We have presented a comprehensive study on text-guided target speech extraction that leverages inter-speaker relative cues across a wide range of speech attributes. By providing a theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative cues versus independent cues on comparable continue-valued attributes.
Experimental results show that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE. Analysis of classification accuracy across cue types indicates that while some cues have strong discriminative power, certain text cues are even more effective than conventional audio-based methods, highlighting their practical potential. Finally, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction. Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of leveraging advanced separation models in TSE frameworks
Overall, our findings underscore the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE. The proposed framework not only mitigates speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.

### v2

@@ IV-D Relative Cues versus Independent Cues
Experimental Setup: We then evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework. For discrete-valued attributes, relative and independent cues provide the same level of discriminability; therefore, we restrict our evaluation to originally continuous-valued attributes. In total, eight cues are considered, of which two—age and temporal order—are excluded. The age cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable. Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (young: 7063; youth: 10310; middle: 2627; senior: 0). The temporal order cue, defined as the difference in onset time between the target and interfering speech, is inherently relational and therefore does not admit a meaningful absolute categorization. Consequently, the evaluation focuses on the remaining six cues: pitch, pitch range, loudness, distance, speaking rate, and speaking duration.
Separated speech samples are selected from the TF-Locoformer separation model based on mixtures associated with these cue types. Two binary classification models are trained: one guided by relative-cue text prompts and the other by independent-cue text prompts.
Independent Cues
  • Pitch: low / normal / high
  • Pitch range: narrow / normal / wide
  • Loudness: quiet / normal / loud
  • Distance: near / far
  • Speaking rate: slow / normal / fast
  • Speaking duration: short / long
Relative Cues
Relative cue categories are defined based on the process explained in Section II and the thresholds specified in Table I:
  • Pitch: lower / similar / higher
  • Pitch range: narrower / similar / wider
  • Loudness: quieter / similar / louder
  • Distance: nearer / similar / farther
  • Speaking rate: slower / similar / faster
  • Speaking duration: shorter / similar / longer
TABLE VI: Metrics for Independent and Relative Cues
@@
Pitch non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
Pitch all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
@@
We use audio mixtures from the dataset described in Section IV that are annotated with the six relative cues—pitch, pitch range, loudness, distance, speaking rate, and speaking duration. We reuse the same predefined templates structure as described in Section II to generate prompts for independent cues. To isolate the effect of cue representation, we control for linguistic variability by regenerating relative-cue prompts using the same predefined templates as those used for independent cues. This ensures that both prompt types share identical sentence structure and phrasing, differing only in whether the cue is expressed in relative or independent terms. Consequently, any performance differences can be attributed specifically to the cue formulation rather than to variations in prompt wording or style. For example, a regenerated relative-cue prompt for a given target speech sample is “Please extract the speaker with a higher pitch in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch in the audio.”
During training, pairs of audio samples and prompts in which both speech signals belong to the same absolute category or the “similar” relative category are excluded. During evaluation, all pairs of audio samples and prompts are retained, with target labels derived from the PIT results of the separation model. Classification accuracy is reported under different conditions. For independent cues, three conditions are considered: (i) same category, (ii) adjacent categories (e.g., slow–normal for speaking rate), and (iii) distinctly different categories (e.g., slow–fast for speaking rate). For relative cues, two conditions are considered: (i) same category (“similar”) and (ii) different categories. We further analyze cases in which target–interference pairs fall into different relative-cue categories but are quantized into the same, adjacent, or distinct absolute categories, highlighting scenarios where independent cues provide limited discriminability while relative cues remain informative.
Results and Analysis: We compare overall classification accuracy across cue types to evaluate the effectiveness of relative versus independent cues for TSE. For independent cues, pairs in the same category are labeled “same,” and those in different categories “non-same.” For relative cues, pairs in the same category are labeled “similar,” and those in different categories “non-similar.” Results are reported over six continuous-valued attributes.
For independent cues, we report classification accuracy along with SI-SDRi and PESQ for the “non-same” and “all” groups. For relative cues, we report these metrics for the “non-similar” and “similar” groups. We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues. Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples. The results are summarized in Table VI.
The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues. Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information. In particular, for pitch, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
@@
For the distance and loudness cues, using relative categories assigns more samples to the “similar” group than absolute categories. This is due to the limited ranges of simulated speaker-to-microphone distances and SIR values. In our dataset, distances range from 0.3 to 1.5 m, with a 0.5 m threshold for distinguishing speakers, leading many pairs to be classified as having similar distances. A similar effect is observed for the loudness cue: loudness is approximated using SIR, which is constrained to the range of -6 to 6 dB in our simulation. With a loudness cue threshold of 3 dB, nearly half of the samples are categorized as having similar loudness.
It is informative to examine the specific categories of target–interference samples. Fig. 5 presents the classification accuracy and sample distribution for both independent (Fig. 5(a)) and relative (Fig. 5(b)) cues, using speaking rate as an example. Compared with the relative speaking rate cue, the independent cue results in more samples being categorized into the “same” group (slow–slow, normal–normal, fast–fast), which cannot be distinguished. In contrast, the relative cue preserves the fine-grained distinctions among these samples, supporting the assumption in Section II. For the independent speaking rate cue, accuracy is highest when target and interference fall into clearly distinct categories (e.g., slow–fast or fast–slow), and decreases for adjacent categories (e.g., slow–normal or normal–fast), as smaller relative differences make classification more difficult.
@@
To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as “non-similar” and examine whether they fall into the same, adjacent, or distinct absolute categories. Using speaking rate as an example (Fig. 6), we report classification accuracy for these conditions on both the full set and the relative-different subset, along with their sample distribution. Accuracy increases with speaking rate difference: in the same condition, it is 48.9% but improves to 79.4% for “non-similar” pairs, indicating added discriminative power from relative comparison. In the adjacent condition, accuracy rises from 87.3% to 89.5%, while in the distinct condition it reaches 98.3% in both cases. These results show that performance strongly correlates with speaking rate difference, with relative cues particularly benefiting cases where absolute quantization obscures pairwise distinctions.
@@ V Conclusion
We have presented a comprehensive study on text-guided target speech extraction that leverages inter-speaker relative cues across a wide range of speech attributes. This is achieved through theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative versus independent cues on comparable continuous-valued attributes.
Experimental results demonstrate that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE. Analysis of classification accuracy across cue types indicates that, while some cues have strong discriminative power, certain textual cues are even more effective than conventional audio-based methods, highlighting their practical potential. Furthermore, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction. Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of incorporating advanced separation models into TSE frameworks.
Overall, our findings highlight the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE. The proposed two-stage framework not only mitigates the speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.